OpenStudy (anonymous):

find the value of x in the expression (2a^4b^2)^x=4a^8b^4 A.x=2 B.x=3 C.x=4 D.x=5 Help please need help

OpenStudy (anonymous):

A

OpenStudy (anonymous):

thanks againg

jimthompson5910 (jim_thompson5910):

(2a^4b^2)^x=4a^8b^4 (2^x)*(a^4)^x*(b^2)^x=4a^8b^4 The coefficients are 2^x (left side) and 4 (right side) So 2^x = 4

jimthompson5910 (jim_thompson5910):

Solve that and you'll get x = 2

OpenStudy (anonymous):

ok thanks